0.858 g of TiCl4 reacts with excess oxygen according to the equation
TiCl4 (s) + O2 (g)  TiO2 (s) + 2Cl2 (g)
What is the percentage yield of TiO2 if 0.287 g of TiO2 was formed?
Molar masses of TiCl4 and TiO2 are, respectively, 189.67 and 79.87 g/mol
1
Expert's answer
2020-11-09T13:49:54-0500

"\begin{aligned} 

TiCl_{4_(s)} + O_{2_(g)} \longrightarrow TiO_{2_(s)} + 2Cl_{2_(g)}

\end{aligned}"

According to the chemical reaction above;

1 mole of "TiCl_4" produces 1mole of "TiO_2"


0.857g of "TiCl_4" = "\frac{0.858}{189.67}mol = 0.0452mol"


"\therefore" 0.00452 mol of "TiCl_4" should produce 0.00452 mol of "TiO_2"

0.00452 mol of "TiO_2" = 0.0452(79.87)g = 0.361g


Therefore 0.361g of "TiO_2" is the theoretical yield;

and 0.287g is the amount of "TiO_2" actually produced


"\begin{aligned}
\% yield = \dfrac{Actual\ yield}{Theoretical\ yield} × 100\%
\end{aligned}" "\begin{aligned}
\\
\% yield = \dfrac{0.287}{0.361} × 100\%
\end{aligned}"

"%yield =\dfrac{Actual\ yield}{Theoretical\ yield}"

"\begin{aligned}
\% yield = 79.5 \%
\end{aligned}"